# GenEDA: Unleashing Generative Reasoning on Netlist via Multimodal Encoder-Decoder Aligned Foundation Model

### v1

@@ V GenEDA Encoder-Decoder Framework
Netlist encoder and LLM decoder in GenEDA. Before discussing the alignment mechanisms, we briefly introduce the models used in GenEDA. For the encoder, we employ NetTAG [19], the state-of-the-art encoder capable of handling post-synthesis netlists, whereas most prior netlist encoders are limited to the AIG format. NetTAG introduces a text-attributed graph format for netlists and employs a multimodal architecture: it encodes the gate function via symbolic Boolean expressions by using an LLM encoder and captures global circuit structure through a graph transformer. This results in rich embeddings that encode both functional and structural information. On the decoder side, we consider both general-purpose LLMs, such as OpenAI GPT and DeepSeek, and circuit-specific LLMs like RTLCoder [20], which is fine-tuned for forward RTL code generation.
@@ V-A Paradigm 1: Embedding-Based Alignment
Paradigm 1 overview. Fig. 4 (a) shows our paradigm 1, which aligns trainable open-source LLM decoders $\mathbf{f}_{\text{dec}}$ with the embeddings generated by the netlist encoder $\mathbf{f}_{\text{enc}}$ from netlist graph (i.e., text-attributed graph) modality. The main challenge is the modality gap between the encoder-generated netlist embeddings and the text embeddings expected by the decoder, as these lie in fundamentally different latent spaces. To address this, our paradigm 1 introduces a modality connector $\mathbf{f}_{\text{c}}$, which acts as a “circuit tokenizer”, transforming the circuit embeddings from our encoder into text-alike embeddings compatible with the LLM decoder. The connector $\mathbf{f}_{\text{c}}$ is cross-modally instruction-tuned with the LLM $\mathbf{f}_{\text{dec}}$, enabling deep embedding-level alignment in the shared latent space. We describe our aligned model architecture and training method below.
Embedding alignment architecture. As shown in Fig. 4 (a), given an input netlist $\mathcal{N}$, we represent it in two modalities: text-attributed graph $\mathcal{G}_{\mathcal{N}}$ and netlist Verilog code $\mathcal{T}_{\mathcal{N}}$ in text format. For $\mathcal{G}_{\mathcal{N}}$, we employ our pre-trained netlist encoder $\mathbf{f}_{\text{enc}}$ to process the netlist to generate the netlist graph-level embedding $N_{\text{g}}$. For the text modality $\mathcal{T}_{\mathcal{N}}$, it is directly tokenized into language embeddings $H_{\text{t}}$ and fed into the LLM decoder $\mathbf{f}_{\text{dec}}$.
@@ V-B Paradigm 2: Prediction-Based Alignment
Paradigm 2 overview. In addition to trainable open-source LLMs, advanced LLMs are often frozen due to commercial or computational limitations, yet they excel in reasoning and support longer input contexts. Unlike paradigm 1, which aligns coarse-grained graph-level embeddings $N_{g}$, our paradigm 2 leverages fine-grained gate-level text and frozen advanced LLMs to support the more challenging exact RTL code reasoning task. As Fig. 4 (b) shows, we align our multimodal encoder $\mathbf{f}_{\text{enc}}$ with frozen LLM decoders $\mathbf{f}_{\text{dec}}$ by using the encoder’s fine-grained gate functional predictions $\mathcal{T}_{\text{pred}}$ as textual inputs for LLMs. These predictions $\mathcal{T}_{\text{pred}}$ provide detailed gate-level analysis of netlist functionalities, which are then used by LLMs to summarize and generate high-level functionality. This paradigm enables seamless integration without modifying the pre-trained frozen decoder. We introduce the details of the task and our alignment method below.
Prediction alignment architecture. To generate functional predictions $\mathcal{T}_{\text{pred}}$ with the encoder $\mathbf{f}_{\text{enc}}$, we fine-tune our pre-trained encoder to enable gate functionality identification with the task proposed in [36]. This task involves classifying gates into five high-level function types defined in their arithmetic RTL code: adder, multiplier, comparator, subtractor, and controller. During fine-tuning, the encoder $\mathbf{f}_{\text{enc}}$ remains frozen, and only the additional function predictor MLP $\mathbf{f}_{\text{p}}$ is trained for gate-type classification using cross-entropy loss, as formulated below.
@@
Chain-of-Thought for RTL code reasoning. Leveraging the annotated netlist, we employ the Chain-of-Thought (CoT) technique to improve the LLM in this challenging exact RTL code reasoning. The CoT prompt decomposes reasoning tasks into two sequential steps: (1) Reason about the word-level arithmetic function description based on the given circuit netlist with gate annotations. (2) Generate RTL code to implement the identified arithmetic function, ensuring the use of word-level operations and avoiding bit-level operations.
@@ VI Experiments
Figure 6: Case study for Task 1. Comparison between the model-generated response and the ground truth specification for a circuit module.
Figure 7: Prompt used in GPT-assisted evaluation.
@@ VI-A Experimental Settings
Circuit dataset preparation. For the circuit encoder and Task 1 & 2, we collect circuit datasets from various open-source RTL code benchmarks, including ITC99 [40], OpenCores [41], Chipyard [42], and VexRiscv [43]. In Task 1 and 2, for large sequential circuits, we split them into multiple sub-circuits, following the techniques in [19]. After splitting, we collect 25k subcircuits, which are augmented by functional equivalent transformation by Yosys [44] to create a total of 50k samples. We randomly sample 400 subcircuits for testing, ensuring that no subcircuits from the same circuit are included in the training set. Since no specification documents are available, we generate functional specifications from the RTL context using GPT-4o as the ground-truth for instruction-tuning. These generated specifications are then verified by human engineers, and those that do not contain valid functionality are excluded. For Task 3, we use the open-source arithmetic RTL code from [36]. Please note that in all tasks, only low-level netlists are provided for reverse reasoning, with high-level specifications and RTL code being unknown to models. All RTL designs are synthesized into netlists using the Synopsys Design Compiler with the NanGate 45nm technology library. We provide detailed statistics of our netlist dataset in TABLE II, including the number of circuits, the average number of text tokens, and the average number of gates.
@@
Figure 8: Case study for Task 3. By utilizing gate function predictions from GenEDA’s encoder, the LLM in GenEDA can reason about the word-level arithmetic functionality of bit-blasted netlists and accurately reconstruct the corresponding RTL code (blue box). In contrast, without these predictions, LLMs (e.g., GPT-4o) generate only bit-level operations, leading to significantly lower reconstruction accuracy (red box).
Model and training. For the encoder model in GenEDA, we employ the state-of-the-art netlist encoder NetTAG [19] as the backbone. For the decoder LLMs in GenEDA, we fine-tune the DeepSeek-Coder [45] 1.3B and 6.7B models in the trainable embedding-based alignment (Paradigm 1), and directly leverage the commercial APIs of OpenAI GPT-4o and DeepSeek-V3 [46] as the frozen LLMs in the prediction-based alignment (Paradigm 2). The training process utilizes DeepSpeed ZeRO and LoRA techniques. In alignment paradigm 1, we adopt a three-layer MLP with dimensions 768, 2048, and 4096 to transform the netlist embedding (768 dimensions) into the LLM word embeddings (4096 dimensions). This connector can be further explored using more advanced methods, such as Q-Former and cross-attention mechanism as in [47, 48, 49] [^2] Complex connectors like Q-Former are typically used with frozen open-source LLMs [48, 49], while simpler connectors such as MLPs are often sufficient when the LLM is trainable [47].. GenEDA is instruction-tuned using LoRA on the full task-specific dataset for one epoch. In alignment paradigm 2, the function predictor MLP for encoder fine-tuning contains three layers with a hidden dimension of 256. Experiments are conducted on a server equipped with 8 NVIDIA A800 80G PCIe GPUs.
@@ VI-C Result of Specification Reasoning (Task 1 & 2)
Baseline solutions. To ensure a comprehensive comparison, we evaluate both general-purpose and circuit-adapted LLMs as baseline methods. For general-purpose LLMs, we use advanced commercial models like GPT-4o and DeepSeek-V3, as well as lightweight open-source models like DeepSeek Coder 1B and 7B. For circuit-adapted LLMs, we select RTLCoder-7B [20], a representative LLM fine-tuned for spec-to-RTL generation. These baselines take netlist text and the question prompt as input and generate the specification text.
Comparison with baselines. TABLE III presents the evaluation results for specification reasoning in Task 1 and Task 2, comparing GenEDA models with baseline models. For both tasks, after aligning with encoder embeddings through multimodal instruction tuning, our GenEDA models (1B and 7B), based on DeepSeek Coder 1B and 7B, significantly outperform all general-purpose LLMs across all textual semantic similarity metrics. Additionally, although RTLCoder [20] is fine-tuned for RTL code generation from specifications, it performs poorly on both tasks, even underperforming its base model, DeepSeek-7B. This is primarily due to the substantial differences between the tasks. Notably, for the GPT scores, which analyze the similarity between generated specification with ground truth with GPT, GenEDA scored much higher than the baseline LLMs.
@@
Natural language specification reasoning case study. We present detailed case studies for these results in Fig. 6. The figure showcases a comparison between a model-generated response and the corresponding ground truth specification for Task 1, function description generation. The model accurately generates key functionality of the specification, aligning closely with the ground truth. For example, in the functionality section, the model effectively describes how the module handles various control and status registers and memory protection configuration, which matches the ground truth’s detailed explanation of register states and operations. These results underline GenEDA’s capability to generate high-level natural language descriptions from low-level netlist inputs.
@@ VI-D Result of RTL Code Reasoning (Task 3)
Notably, even state-of-the-art commercial LLMs like GPT-4o and DeepSeek-V3 achieve less than a 20% functional success rate on our benchmark, highlighting the difficulty of this task. In contrast, existing specification-to-RTL generation benchmarks like RTLLM [38] and VerilogEval [39] report over 60% success with off-the-shelf LLMs like GPT-4, emphasizing the challenge of reverse reconstruct high-level RTL code from low-level bit-blasted netlists.
@@ VI-E Ablation Study
Effectiveness of alignment with encoders. GenEDA alignment paradigm 1 is achieved by cross-modally fine-tuning the LLMs using encoder embeddings. In this ablation study, we remove the encoder alignment and only perform supervised fine-tuning (SFT) of the LLMs using the same prompts and labels as in GenEDA. As shown in Fig. 9 (a), removing the encoder alignment significantly decreases model performance on Task 1 and Task 2 across all metrics. Notably, the GPT scores drop sharply from 0.62 to 0.32 on Task 1 and from 0.51 to 0.21 on Task 2, highlighting the effectiveness of embedding alignment.
Impact of encoder quality. GenEDA alignment paradigm 2 heavily relies on the accuracy of gate functionality classification from the encoder. In this ablation study, we replace the high-quality encoder NetTAG [19] with a less effective baseline, GNN-RE [36]. This results in a significant drop in classification accuracy from 97% to 83%. Consequently, the performance of reconstructed RTL code also degrades, with syntax accuracy dropping from 96% to 91% and functional accuracy from 69% to only 40%, as shown in Fig. 9 (b). This demonstrates the critical importance of using high-quality encoders in generative tasks that involve netlists.
@@ VII-A Extending GenEDA Alignment to Other Circuit Stages and Tasks
Beyond the netlist stage addressed in this work, GenEDA’s encoder-decoder alignment framework can be extended to various stages in the circuit design flow. At the RTL stage, even though the same RTL functionality might be present, different circuit structures can yield significantly varying PPA characteristics. By leveraging structural RTL information captured by RTL encoders, LLMs can enable structure-aware generation, potentially generating more optimized RTL code with better PPA characteristics. Additionally, at the layout stage, GenEDA can be adapted to handle cross-modal inputs, such as layout images and netlist graphs. This might enable the direct generation of macro positions on a chip by learning from image representations, thus enabling more efficient and optimized physical design generation.
VII-B Potential Application of Netlist Generative Reasoning
@@ VIII Conclusion and Future Work
In this paper, we present GenEDA, a framework that aligns multimodal circuit encoders and decoders for advanced generative circuit functional reasoning tasks. We align the state-of-the-art netlist encoder with both trainable and frozen LLM decoders through two alignment paradigms. Our experiments and ablation studies demonstrate the effectiveness of this approach, with GenEDA significantly enhancing the performance of state-of-the-art LLMs, showcasing the critical role of integrating both graph and text circuit modalities for complex netlist tasks. Future work will explore extending this alignment framework to other stages of the circuit design flow, such as RTL code generation and layout-stage tasks, further enhancing the capabilities of GenEDA for diverse EDA applications.

### v2

@@ V GenEDA Encoder-Decoder Framework
Netlist encoder and LLM decoder in GenEDA. Before discussing the alignment mechanisms, we briefly introduce the models used in GenEDA. For the encoder, we employ NetTAG [19], the state-of-the-art encoder capable of handling post-synthesis netlists, whereas most prior netlist encoders are limited to the AIG format. NetTAG introduces a text-attributed graph format for netlists and employs a multimodal architecture: it encodes the gate function via symbolic Boolean expressions using an LLM encoder and captures global circuit structure through a graph transformer. This results in rich embeddings that encode both functional and structural information. On the decoder side, we consider both general-purpose LLMs, such as OpenAI GPT and DeepSeek, and circuit-specific LLMs like RTLCoder [20], which is fine-tuned for forward RTL code generation.
@@ V-A Paradigm 1: Embedding-Based Alignment
Paradigm 1 overview. Fig. 4 (a) shows our paradigm 1, which aligns trainable open-source LLM decoders $\mathbf{f}_{\text{dec}}$ with the embeddings generated by the netlist encoder $\mathbf{f}_{\text{enc}}$ from the netlist graph (i.e., text-attributed graph) modality. The main challenge is the modality gap between the encoder-generated netlist embeddings and the text embeddings expected by the decoder, as they lie in fundamentally different latent spaces. To address this, our paradigm 1 introduces a modality connector $\mathbf{f}_{\text{c}}$, which acts as a “circuit tokenizer”, transforming the circuit embeddings from our encoder into text-alike embeddings compatible with the LLM decoder. The connector $\mathbf{f}_{\text{c}}$ is cross-modally instruction-tuned with the LLM $\mathbf{f}_{\text{dec}}$, enabling deep embedding-level alignment in the shared latent space. We describe our aligned model architecture and training method below.
Embedding alignment architecture. As shown in Fig. 4 (a), given an input netlist $\mathcal{N}$, we represent it in two modalities: text-attributed graph $\mathcal{G}_{\mathcal{N}}$ and netlist Verilog code $\mathcal{T}_{\mathcal{N}}$ in text format. For $\mathcal{G}_{\mathcal{N}}$, we employ our pre-trained netlist encoder $\mathbf{f}_{\text{enc}}$ to process the netlist and generate the netlist graph-level embedding $N_{\text{g}}$, which contains the structural and functional netlist information. For the netlist Verilog text $\mathcal{T}_{\mathcal{N}}$, we keep only the signal names and leave the structural details to be captured by the encoder model. As shown in Fig. 5, the signal name text is combined with the question prompt instruction to form the textual input $\mathcal{T}_{\text{in}}$. This input $\mathcal{T}_{\text{in}}$ is then tokenized into language embeddings $H_{\text{t}}$ and fed into the LLM decoder $\mathbf{f}_{\text{dec}}$.
@@ V-B Paradigm 2: Prediction-Based Alignment
Paradigm 2 overview. In addition to trainable open-source LLMs, advanced LLMs are often frozen due to commercial or computational limitations, yet they excel in reasoning and support longer input contexts. Unlike paradigm 1, which aligns coarse-grained graph-level embeddings $N_{g}$, our paradigm 2 leverages fine-grained gate-level text and frozen advanced LLMs to support the more challenging exact RTL code reasoning task. As Fig. 4 (b) shows, we align our multimodal encoder $\mathbf{f}_{\text{enc}}$ with frozen LLM decoders $\mathbf{f}_{\text{dec}}$ by using the encoder’s fine-grained gate functional predictions $\mathcal{T}_{\text{pred}}$ as textual inputs for LLMs. These predictions $\mathcal{T}_{\text{pred}}$ provide detailed gate-level analysis of netlist functionalities, which are then used by LLMs to summarize and generate high-level functionality. This paradigm enables seamless integration without modifying the pre-trained frozen decoder. Below, we introduce the details of the task and our alignment method.
Prediction alignment architecture. To generate functional predictions $\mathcal{T}_{\text{pred}}$ with the encoder $\mathbf{f}_{\text{enc}}$, we fine-tune our pre-trained encoder to enable gate functionality identification with the task proposed in [37]. This task involves classifying gates into 5 high-level function types defined in their arithmetic RTL code, including adder, multiplier, comparator, subtractor, and controller. During fine-tuning, the encoder $\mathbf{f}_{\text{enc}}$ remains frozen, and only the additional function predictor MLP $\mathbf{f}_{\text{p}}$ is trained for gate function classification using cross-entropy loss, as formulated below.
@@
Chain-of-Thought for RTL code reasoning. Leveraging the annotated netlist, we employ the Chain-of-Thought (CoT) technique to improve the LLM in this challenging exact RTL code reasoning. The CoT prompt decomposes reasoning tasks into two sequential steps: (1) Reason about the word-level arithmetic function description based on the given circuit netlist with gate annotations. (2) Generate RTL code to implement the identified arithmetic function, ensuring word-level RTL operations and avoiding bit-level operations.
@@ VI Experiments
  • RTLCoder [20] is fine-tuned for generating RTL code from specifications. In our setting, we adopt the fine-tuned model as a circuit-specific LLM baseline to generate specifications from netlists reversely.
Figure 6: Case study for Task 1. Comparison between the response from GenEDA and the ground truth specification for a circuit module.
Figure 7: Prompt used in GPT-assisted evaluation (i.e., GPT Score).
@@ VI-A Experimental Settings
Circuit dataset preparation. For the circuit encoder and Task 1 & 2, we collect circuit datasets from various open-source RTL code benchmarks, including ITC99 [41], OpenCores [42], Chipyard [43], and VexRiscv [44]. In Task 1 and 2, for large sequential circuits, we split them into multiple sub-circuits, following the techniques in [19]. After splitting, we collect 25k subcircuits, which are augmented by functional equivalent transformation by Yosys [45] to create a total of 50k samples. We randomly sample 400 subcircuits for testing, ensuring that no subcircuits from the same circuit are included in the training set. Since no specification documents are available, we generate functional specifications from the RTL context using GPT-4o as the ground-truth for instruction-tuning. Human engineers then verify these generated specifications, and those that do not contain valid functionality are excluded. For Task 3, we use the open-source arithmetic RTL code from [37]. Please note that in all tasks, only low-level netlists are provided for reverse reasoning, with high-level specifications and RTL code being unknown to models. All RTL designs are synthesized into netlists using the Synopsys Design Compiler with the NanGate 45nm technology library. We provide detailed statistics of our netlist dataset in TABLE II, including the number of circuits, the average number of text tokens, and the average number of gates.
@@
Figure 8: Case study for Task 3. By utilizing gate function predictions from GenEDA’s encoder, the LLM in GenEDA can reason about the word-level arithmetic functionality of bit-blasted netlists and accurately reconstruct the corresponding RTL code (blue box). In contrast, without these predictions, LLMs (e.g., GPT-4o) generate only bit-level operations (red box), leading to significantly lower reconstruction accuracy.
Model and training. For the encoder model in GenEDA, we employ the state-of-the-art netlist encoder NetTAG [19] as the backbone. For the decoder LLMs in GenEDA, we fine-tune the DeepSeek-Coder [46] 1B and 7B models in the trainable embedding-based alignment (Paradigm 1), and directly leverage the commercial APIs of OpenAI GPT-4o and DeepSeek-V3 [47] as the frozen LLMs in the prediction-based alignment (Paradigm 2). The training process utilizes DeepSpeed ZeRO and LoRA techniques. In alignment paradigm 1, we adopt a three-layer MLP with dimensions 768, 2048, and 4096 to transform the netlist embedding (768 dimensions) into the LLM word embeddings (4096 dimensions). This connector can be further explored using more advanced methods, such as Q-Former and cross-attention mechanism as in [48, 49, 50] [^3] Complex connectors like Q-Former are typically used with frozen open-source LLMs [49, 50], while simpler connectors such as MLPs are often suitable when the LLM is trainable [48].. GenEDA is instruction-tuned using LoRA on the full task-specific dataset for one epoch. In alignment paradigm 2, the function predictor MLP for encoder fine-tuning contains three layers with a hidden dimension of 256. Experiments are conducted on a server equipped with 8 NVIDIA A800 80G PCIe GPUs.
@@ VI-C Result of Specification Reasoning (Task 1 & 2)
Baseline solutions. To ensure a comprehensive comparison, we evaluate both general-purpose and circuit-adapted LLMs as baseline methods. For general-purpose LLMs, we use advanced commercial models like GPT-4o and DeepSeek-V3, as well as lightweight open-source models like DeepSeek-Coder 1B and 7B. For circuit-adapted LLMs, we select RTLCoder-7B [20], a representative LLM fine-tuned for spec-to-RTL generation. These baselines take netlist text and the question prompt as input and generate the specification text.
Comparison with baselines. TABLE III presents the evaluation results for specification reasoning in Task 1 and Task 2, comparing GenEDA models with baseline models. For both tasks, after aligning with netlist encoder embeddings through multimodal instruction tuning, our GenEDA models (1B and 7B), based on DeepSeek Coder 1B and 7B, significantly outperform all general-purpose LLMs across all textual semantic similarity metrics. Additionally, although RTLCoder [20] is fine-tuned for RTL code generation from specifications, it performs poorly on both tasks, even underperforming its base model, DeepSeek-7B. This is primarily due to the substantial differences between the tasks. Notably, for the GPT scores, which analyze the similarity between generated specifications and ground truth with GPT, GenEDA scored much higher than the baseline LLMs.
@@
Natural language specification reasoning case study. We present detailed case studies for these results in Fig. 6. We compare a model-generated response and the corresponding ground truth specification for Task 1, function description generation. The model accurately generates key functionality of the specification, aligning closely with the ground truth. For example, in the functionality section, the model effectively describes how the module handles various control and status registers and memory protection configuration, which matches the ground truth’s detailed explanation of register states and operations. These results underline GenEDA’s capability to generate high-level natural language descriptions from low-level netlist inputs.
@@ VI-D Result of RTL Code Reasoning (Task 3)
Notably, even state-of-the-art commercial LLMs like GPT-4o and DeepSeek-V3 achieve less than a 20% functional success rate on our benchmark, highlighting the difficulty of this task. In contrast, existing specification-to-RTL generation benchmarks like RTLLM [39] and VerilogEval [40] report over 60% success with off-the-shelf LLMs like GPT-4, emphasizing the challenge of reverse reconstructing high-level RTL code from low-level bit-blasted netlists.
@@ VI-E Ablation Study
Effectiveness of alignment with encoders. GenEDA alignment paradigm 1 is achieved by cross-modally fine-tuning the LLMs using netlist encoder embeddings. In this ablation study, we remove the encoder alignment and only perform supervised fine-tuning (SFT) of the LLMs using the same prompts and labels as in GenEDA. As shown in Fig. 9 (a), removing the encoder alignment significantly decreases model performance on Task 1 and Task 2 across all metrics. Notably, the GPT scores drop sharply from 0.62 to 0.32 on Task 1 and from 0.51 to 0.21 on Task 2, highlighting the effectiveness of embedding alignment.
Impact of encoder quality. GenEDA alignment paradigm 2 heavily relies on the accuracy of gate functionality classification from the encoder. In this ablation study, we replace the high-quality encoder NetTAG [19] with a less effective baseline, GNN-RE [37]. This results in a significant drop in classification accuracy from 97% to 83%. Consequently, the performance of reconstructed RTL code also degrades, with syntax accuracy dropping from 96% to 91% and function accuracy decreasing from 69% to only 40%, as shown in Fig. 9 (b). This demonstrates the importance of using high-quality encoders and their impact on the generation after alignment.
@@ VII-A Extending GenEDA Alignment to Other Circuit Stages and Tasks
Beyond the netlist stage addressed in this work, GenEDA’s encoder-decoder alignment framework can be extended to various stages in the circuit design flow. At the RTL stage, even though the same RTL functionality can be generated, different circuit structures can yield significantly varying PPA characteristics. By leveraging structural RTL information captured by RTL encoders, LLMs can potentially enable structure-aware generation, generating more optimized RTL code with better PPA metrics. Additionally, at the layout stage, GenEDA can be adapted to handle cross-modal inputs, such as layout images and netlist graphs. This might enable the direct generation of macro positions on a chip by learning from image representations, thus enabling more efficient and optimized physical design generation.
VII-B Potential Application of Generative Netlist Functional Reasoning
@@ VIII Conclusion and Future Work
In this paper, we present GenEDA, a framework that aligns multimodal circuit encoders and decoders for advanced generative netlist functional reasoning tasks. We align the state-of-the-art netlist encoder with both trainable and frozen LLM decoders through two alignment paradigms. Our experiments and ablation studies demonstrate the effectiveness of this approach, with GenEDA significantly enhancing the performance of state-of-the-art LLMs, showcasing the critical role of integrating both graph and text circuit modalities for complex netlist tasks. Future work will explore extending this alignment framework to other stages of the circuit design flow, such as RTL code generation and layout-stage tasks, further enhancing the capabilities of GenEDA for diverse EDA applications.
IX Acknowledgement
This research was supported by Hong Kong Research Grants Council (RGC) CRF-YCRG Grant C6003-24Y, GRF 16200724, and ACCESS – AI Chip Center for Emerging Smart Systems, sponsored by the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government.